Concord Company produces one product, a putter called GO-Putter. Concord uses a standard cost system and determines that it should take one hour of direct labor to produce one GO-Putter. The normal production capacity for this putter is 120,000 units per year. The total budgeted overhead at normal capacity is $1,080,000 comprised of $420,000 of variable costs and $660,000 of fixed costs. Concord applies overhead on the basis of direct labor hours. During the current year, Concord produced 75,300 putters, worked 87,600 direct labor hours, and incurred variable overhead costs of $169,425 and fixed overhead costs of $624,700. Compute the predetermined variable overhead rate and the predetermined fixed overhead rate. (Round answers to 2 decimal places, e.g. 2.75.) Variable Fixed Predetermined Overhead Rate $ $ Compute the applied overhead for Concord for the year. Overhead Applied $ Compute the total overhead variance. Total Overhead Variance $
Answer #1

a. The predetermined variable overhead rate = $420,000 /120,000 = $3.50 per DLH

The predetermined fixed overhead rate = $660,000 /120,000 = $5.50 per DLH

b. Standard Direct Labor for manufacturing 75300 units = 75300*1 = 75300 Hour

The total predetermined overhead rate for the year = 3.50 + 5.50 = $9

The applied total overhead for the year = 75300 * 9 = $677,700

c. Total Overhead Variance = Actual Overhead - Standard Overhead applied

Total Overhead Variance = 169425+624700-677700 = $116,425 Unfavorable